OpenStudy (anonymous):

Which equation represents the line passing through the points (3,2) and (-9,6)? x-3y=9 x+3y=9 3x-y=-9 3x+y=9

OpenStudy (anonymous):

first find the slope using the two point y2-y1/x2-x1 could you do that?

OpenStudy (anonymous):

y=4 and x=-12

OpenStudy (anonymous):

yeah divide them

OpenStudy (anonymous):

that is -0.3 repeating

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

simple -1/3 is your slope

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what's next

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Never mind. I got to gp

OpenStudy (anonymous):

Thanks anyway

OpenStudy (anonymous):

now use y-y1[=-1/3(x-x1) where y1=2 and x1= 3 put these value an you willl get x+3y=9